OpenStudy (anonymous):

Given the function f(x) = 5(x+4) − 6, solve for the inverse function when x = 19. 1 68 72 84

OpenStudy (anonymous):

@Here_to_Help15

OpenStudy (perl):

first simplify f(x) y = 5x + 20 - 6 y = 5x + 14 now switch x and y , solve for y

OpenStudy (anonymous):

so it be 5x=y+4

OpenStudy (anonymous):

No.

OpenStudy (anonymous):

im confused

OpenStudy (anonymous):

Just interchange the letters for y=5x+14 which should give you x=5y+14

OpenStudy (anonymous):

ahhh

OpenStudy (jdoe0001):

\(\bf f(x) = 5(x+4) - 6\implies y=5x+20-6 \\ \quad \\ {\color{brown}{ y}}=5{\color{blue}{ x}}+14\qquad inverse\implies {\color{blue}{ x}}=5{\color{brown}{ y}}+14 \) and solve for "y"

OpenStudy (anonymous):

alright let me see if i get this

OpenStudy (anonymous):

nopee still confused

OpenStudy (anonymous):

@jim_thompson5910 can u also help

OpenStudy (anonymous):

by finding y, it will give us the inverse of f(x).

OpenStudy (anonymous):

x=5y+14 let us make y the subject of the formula, so subtract 14 from both sides

OpenStudy (anonymous):

would you put the 14 under the 5x

OpenStudy (anonymous):

go ahead and draw what you have so far.

OpenStudy (anonymous):